The school that DeShawn goes to is selling tickets to a fall musical. On the first day of ticket
sales the school sold 5 senior citizen tickets and 6 child tickets for a total of $48. The school took
in $54 on the second day by selling 5 senior citizen tickets and 8 child tickets. What is the price
each of one senior citizen ticket and one child ticket?

Answer by ikleyn(52813)

.

            This problem is to be solved  MENTALLY,  without using equations,  but with the goal to train your mind.


After reading the problem two times, an attentive reader will see momentarily that


    the difference in the bought tickets is 2 child tickets, while the difference in the cost is 54-48 = 6 dollars.


Hence, one child ticket price is 6/2 = 3 dollars.    ANSWER


Then 5 senior tickets cost  48-6*3 = 30 dollars;  hence, one senior ticket price is  30/6 = 5 dollars.    ANSWER